The length of one side of a square tile is 1 2/3 feet. what is the perimeter of the tile in inches?
Answer by ReadingBoosters(3246)

1 ft = 12 in
Convert 1 2/3 ft into a mixed fraction,
(3*1)+2 = 5 so 5/3
5/3 ft x 12 in/ft = 60/3 in or 20 in
As a square, all sides are equal and perimeter is 2l+2w or 2(l+w)
2(20in + 20in)
2(40in)
perimeter is 80inches